OpenStudy (anonymous):

find a general form equation for the line through the pair of points (-7,-2) and (1,6)

OpenStudy (anonymous):

The formula for an equation is \[y = y _{1} + m(x - x _{1})\]where \[m = \frac{ y _{2} - y _{1}}{ x _{2} - x _{1}}\] and \[(x _{1} , y _{1}) ,(x _{2} , y _{2})\] are two points on the line.

OpenStudy (anonymous):

so i need to find the slope

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Then, put it into the formula

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

Yes, then express it in the form of \[ax + by = c\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@TURITW

OpenStudy (anonymous):

could you work it out

OpenStudy (anonymous):

\[y = -2 + \frac{ 6 - -2 }{ 1 - -7 }(x - -7)\]\[y = -2 + x + 7\]\[y = x + 5\]\[x - y = -5\]

OpenStudy (anonymous):

oh ok thanks